# Multi-layer diffusion model of photovoltaic installations

### v2

@@ Abstract
Nowadays, harmful effects of climate change are becoming increasingly apparent. A vital issue that must be addressed is the generation of energy from non-renewable and often polluting sources. For this reason, the development of renewable energy sources is of great importance. Unfortunately, too rapid spread of renewables can disrupt stability of the power system and lead to energy blackouts. One should not simply support it, without ensuring sustainability and understanding of the diffusion process. In this research, we propose a new agent-based model of diffusion of photovoltaic panels. It is an extension of the $q$-voter model that utilizes multi-layer network structure. The novelty is studying both opinion dynamics and diffusion of innovation simultaneously, on a multidimensional structure. The model is analyzed with Monte Carlo simulations and mean-field approximation. The impact of parameters and specifications on the basic properties of the model is discussed. Firstly, we show that for a certain range of parameters, innovation always succeeds, regardless of the initial conditions. Secondly, that the mean-field approximation gives qualitatively the same results as computer simulations, even though it does not utilize knowledge of the network structure.
@@ 1 Introduction
Monte Carlo (MC) computer simulations are the main research method for ABMs. However, in some limited cases, as for networks being complete or random graphs with a low clustering coefficient, analytical methods such as mean-field (MFA) or pair approximation (PA) can be used [36, 37, 38]. Unfortunately, real world networks are characterized by a high clustering coefficient [18], so often the analytical results obtained this way differ quantitatively from ones given by Monte Carlo simulations [37]. The need to sweep a multidimensional space of input parameters and perform multiple independent repetitions to obtain good statistical parameters makes Monte Carlo simulations very time-consuming. For this reason, we utilize both computer simulations and analytical methods in this research.
@@
A popular agent-based model that already has found its use in studying diffusion of eco-innovation is the $q$-voter model [20, 23, 39]. In this model, conformity is the basic form of social response. Agents are characterized by a single binary variable denoting their opinion and placed in nodes of some underlying graph structure. Their opinions change upon impact with so-called groups of influence, i.e. $q$ of the agent’s neighbors chosen randomly, but only if a given group presents an unanimous opinion. A well-established extension to the $q$-voter model is the addition of independence – a probability that an agent acts and changes opinion independently of the group of influence [40]. The $q$-voter model has been already examined on multi-layer networks, but to the best of our knowledge only ones consisting of two identical layers [35, 41].
@@
In this paper, we introduce a new model based on the aforementioned $q$-voter model with independence. Our goal is to capture both opinion dynamics and diffusion of innovation jointly. Hence, we introduce (compared to the original $q$-voter model) a second agent’s attribute, in addition to the opinion – an adoption state. Similarly to the opinion, it is a binary one. Positive value means that an agent possesses a PV installation, negative – it does not. An agent (in this case a household) can acquire knowledge of PVs from two sources. Either it sees panels on the roofs of neighboring households, or communicates with friends/colleagues. To model this twofold dynamic, we utilize a two-layer network as the underlying topology. The first layer of the network depicts the spacial location of an agent, similarly to [42, 43]. From a mathematical point of view it is just a square lattice (SL) with Moore’s neighborhood [44]. At this level, each agent possesses only knowledge of the adoption states of the closest neighborhood, and these adoption states shape agent’s opinion, according to the $q$-voter rule. The agent cannot share opinion or be subject to the opinions of the neighborhood. This is the visual observation part. A single-family house owner sees only PV panels, or lack thereof, on the roofs of neighboring houses, and does not exchange information or opinions with neighbors. The second layer of the network corresponds to structure of contacts and relationships of agents. For this reason, here we use two-dimensional Watts-Strogatz graph (WS2D) [45]. It possesses some characteristics of real world social networks [18, 46] and, in a specific case (when randomness equals 00), reduces to a square lattice. At this level, a connection between two agents implies their familiarity, through which they can exchange opinions, but do not observe their adoption states. In this case, social influence is also described by the $q$-voter model. Lastly, it is an agent’s opinion solely that impacts its adoption state. An agent with a positive opinion can install photovoltaic panels with probability $a_{1}$, one with a negative one can resign with probability $a_{2}$. The reason for this research is to thoroughly examine this mathematical model, to understand how its parameters shape the outcome.
@@
The rest of this article is organized as follows. In the next section, we provide a more detailed description of the model and the methods used. Then, in Section 3, we examine the two variants, AND and OR rules, which were explained above. Both by means of computer simulations, numerical and analytical methods. Finally, in Section 4, we wrap up. Additionally, in Appendix, we consider various initial conditions.
@@ 2.1 Simulation Model
We consider a set of $N$ agents, each of which is characterized by 2 binary variables: adoption $A_{i}=\pm 1$ (adopted or not adopted), and opinion $S_{i}=\pm 1$ (positive or negative), for $i=1,2,\dots,N$. Agents are located on a two-layer network (Fig. 2). The first layer of the network depicts the spacial location of an agent and is represented by a square lattice (SL) with Moore’s neighborhood, meaning that each agent possesses 8 neighbors surrounding it (except for agents in corners and along edges, which have 3 and 5 neighbors, respectively) [44]. At this level, an agent sees only the adoption states of its neighbors. The second layer corresponds to the structure of social ties of agents, which is described by a two-dimensional Watts-Strogatz graph [45]. Here, an agent sees only the opinions of its neighbors. Each layer is a connected simple graph [47], as described in Definition 1.
@@
Let $G=\left[G_{X,i,j}\right]$ denote an adjacency matrix for the two-layer network, with each layer $X$, $X\in\{1,2\}$, being a connected simple graph, i.e. an unweighted, undirected graph containing no graph loops or multiple edges [47]. Then, for $i,j=1,2,\dots,N$:
@@
We investigate the model through Monte Carlo simulations with a random sequential updating scheme. Within a single simulation, the time is measured in so-called Monte Carlo sweeps (MCS) [48]. One MCS consists of $N$ elementary events, each of length $\Delta t=\frac{1}{N}$. In each elementary event, an agent $i$ is chosen at random (uniformly from all $N$). Then, with probability $p$, agent $i$ acts independently and changes its opinion $S_{i}$ randomly. With complementary probability $1-p$, the agent is susceptible to social influence (based on the $q$-voter model [39]), combined from both layers of the network with respect to the variant, AND or OR. Lastly, if agent $i$ has positive opinion $S_{i}=+1$, but negative adoption state $A_{i}=-1$, it changes the latter to positive, $A_{i}\to+1$, with probability $a_{1}$. Otherwise, if it possesses negative opinion $S_{i}=-1$, but positive adoption state $A_{i}=+1$, it looses the latter, i.e. $A_{i}\to-1$, with probability $a_{2}$. Simulation details are shown in Algorithm 1, and the graphical representation of the model in Fig. 2. In Algorithm 1:
@@
    $G$ is an adjacency matrix for the two-layer network, as per Definition 1,
@@
To clarify, we do the following. At the beginning of each independent simulation (trajectory), we set the initial conditions. In simulations, we always start from a fully unadopted, negative state, i.e., $\forall_{i}\,\,A_{i}(0)=-1,\,\,S_{i}(0)=-1$. Then, we perform Monte Carlo sweeps until a determined time horizon $T$ is reached ($T$ is the number of MCS). We repeat independent simulations multiple times with the same set of parameter values for a better statistical accuracy. We also run separate simulations for different sets of parameter values.
@@
$c_{A}=\frac{1}{N}\sum_{i=1}^{N}\left(2A_{i}-1\right),\,\,\,\,c_{S}=\frac{1}{N}\sum_{i=1}^{N}\left(2S_{i}-1\right).$ (2)
@@ 2.2 Mean-Field Approximation
Let $c_{A}$ and $c_{S}$ denote the concentrations of positive adoption states and positive opinions respectively (as per Definition 2). Under the assumption that each layer of the network is a complete graph of size $N\to\infty$, the dynamics of the system is described in the AND variant by:
@@
Under the assumption that each layer of the network is a complete graph of size $N\to\infty$, we can write down probability $\gamma_{A}^{+}$ that a number of agents with positive adoption states will increase by $1$ and probability $\gamma_{A}^{-}$ that it will decrease by $1$:
@@
which gives us Eq. (3), after replacing $a_{2}$ with $ha_{1}$ (Eq. (1)). Eq. (3) is common for both AND and OR variants of the model. Similarly, we can write down probabilities $\gamma_{S}^{+}$ and $\gamma_{S}^{-}$ that a number of agents with positive opinion will increase or decrease by $1$ respectively. These are:
@@
Due to the power of $q$ in Eqs. (4)-(5), time trajectories of the system given by Eqs. (3)-(5) cannot be determined analytically. Hence, one must resort to numerical methods, which we show in Section 3. However, stationary states can be found analytically and there existence can be proven.
@@
Let $c_{A}$ and $c_{S}$ denote the concentrations of positive adoption states and positive opinions respectively (as per Definition 2). Under the assumption that each layer of the network is a complete graph of size $N\to\infty$, for any $q\in\mathbb{N}$, $q\geq 2$, $p\in[0,1]$, $h\in(0,1)$ and $a_{1}\in(0,1]$, there always exists at least one stationary state $(c_{A}^{\prime},c_{S}^{\prime})$, and all stationary states $(c_{A}^{\prime},c_{S}^{\prime})$ must satisfy:
@@
A stationary state $(c_{A}^{\prime},c_{S}^{\prime})$ must satisfy:
@@
When $c_{S}\in\left[\frac{1}{2},1\right]$, $f_{4}(c_{S})$ is a rational function with positive denominator, and hence continuous within that interval. As it is continuous in $\left[\frac{1}{2},1\right]$, $f_{4}(1)=0$ and $f_{4}\left(\frac{1}{2}\right)=1$, it achieves any value of between 00 and $1$. Hence, for any $p\in[0,1]$ there always exists at least one $c_{S}\in[0,1]$ that satisfies Eq. (30). Therefore, in the AND variant, there always exists at least one stationary state.
@@
where
$f_{3}(c_{S})={}$ $\left(c_{S}-\frac{1}{2}\right)(c_{S}+h-c_{S}h)^{q},$
@@
When $c_{S}\in\left[\frac{1}{2},1\right]$, $f_{6}(c_{S})$ is a rational function with positive denominator, and hence continuous within that interval. As it is continuous in $\left[\frac{1}{2},1\right]$, $f_{6}(1)=0$ and $f_{6}\left(\frac{1}{2}\right)=1$, it achieves any value of between 00 and $1$. Hence, for any $p\in[0,1]$ there always exists at least one $c_{S}\in[0,1]$ that satisfies Eq. (36). Therefore, in the OR variant, there always exists at least one stationary state. ∎
@@
First, we show that for $c_{S}\in\left[\frac{1}{2},1\right)$:
@@ 3.1 Simulation Details
First, we examine our model with Monte Carlo computer simulations. In order to measure the outcomes at a macroscopic level, we use concentrations, i.e. fractions, of positive adoption states $c_{A}$ and positive opinions $c_{S}$, see Definition 2. In all the simulations, we study a system of size $N=2500$ (number of agents, which may be interpreted as a borough in a city). As mentioned in the Introduction, the first layer of the two-layer network is always a Square Lattice with Moore’s neighborhood [44], and no periodic boundary conditions, which means that almost every agent has exactly 8 neighbors (except those on the edges of the lattice). We denote it as SL($N=2500$, $m=1$), where $m=1$ indicates the range of a neighborhood on a square grid. The second layer is always a two-dimensional Watts-Strogatz graph [45], with Moore’s neighborhood before rewiring and rewiring probability $\beta=0.2$, and again, without periodic boundary conditions. We wanted a layer that possesses characteristics of a real world social network, but still bears a resemblance to the other layer. We denote it as WS2D($N=2500$, $m=1$, $\beta=0.2$). We study the behavior of the system with respect to the parameters $p$, $a_{1}$ and $a_{2}$. Although the size of the group of influence $q$ is a parameter as well, here we keep it constant, as its impact in the $q$-voter model with independence is well studied [40]. We choose $q=4$ motivated by real life experiments [50]. This will help us to reduce computational efforts slightly. In all the simulations, we start from a fully unadopted, negative state, i.e., $c_{A}(0)=0$, $c_{S}(0)=0$. Lastly, for illustrative purposes, the time horizon $T$ for each simulation is limited to $5000$ Monte Carlo steps (MCS), and the number of time trajectories to $10$.
@@ 3.2 Simulation Model
Figure 3: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND variant (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Let us begin with Fig. 3. It shows evolution of the system, for both variants (AND and OR), and a range of values of $p$, $a_{1}$ and $a_{2}$, where $a_{2}=0.5a_{1}$ in all cases. We can observe that several final outcomes are possible. When $p$ is low, we retain an unadopted state ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$). None of the agents have a positive opinion or adoption state, except for a few rebels. Then, there is a critical value of $p$, above which the system becomes adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$). If the value of $p$ is high, independence surpasses conformity and we end up with a disordered system ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$). There is a visible difference between the two variants: AND and OR. In the first, much lower values of $p$ are required to enter adopted or disordered state than in the latter. Finally, the value of $a_{1}$ itself has no impact on the final state (except for $a_{1}=0$, which is not shown here), only on the time needed to reach it, with the higher value speeding up the process.
Figure 4: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND variant (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.25a_{1}$ in all cases.
What affects the final state, however, is the relationship between $a_{1}$ and $a_{2}$, namely the coefficient $h$ (Eq. (1)). Let us move to Fig. 4. There, $a_{2}$ is twice as small relative to $a_{1}$ as before, $a_{2}=0.25a_{1}$. With this decrease in $a_{2}$ much lower values of $p$ are needed, in both variants, for the system to become adopted. This itself is a rather trivial conclusion, as now the probability of loosing adoption is four times smaller that the one of adopting. There is, however, a secondary effect to it, due to the fact how the first layer of the network impacts opinions in our model. Positive adoption states support positive opinions (and vice versa). For this reason, the adopted state is not only easier to reach, but also more difficult to disorder.

### v3

@@ Abstract
Nowadays, harmful effects of climate change are becoming increasingly apparent. A vital issue that must be addressed is the generation of energy from non-renewable and often polluting sources. For this reason, the development of renewable energy sources is of great importance. Unfortunately, too rapid spread of renewables can disrupt stability of the power system and lead to energy blackouts. One should not simply support it, without ensuring sustainability and understanding of the diffusion process. In this research, we propose a new agent-based model of diffusion of photovoltaic panels. It is an extension of the $q$-voter model that utilizes a multi-layer network structure. The novelty is that both opinion dynamics and diffusion of innovation are studied simultaneously on a multidimensional structure. The model is analyzed using Monte Carlo simulations and the mean-field approximation. The impact of parameters and specifications on the basic properties of the model is discussed. Firstly, we show that for a certain range of parameters, innovation always succeeds, regardless of the initial conditions. Secondly, that the mean-field approximation gives qualitatively the same results as computer simulations, even though it does not utilize knowledge of the network structure.
@@ 1 Introduction
Monte Carlo (MC) computer simulations are the main research method for ABMs. However, in some limited cases, as for networks being complete or random graphs with a low clustering coefficient, analytical methods such as mean-field (MFA) or pair approximation (PA) can be used [36, 37, 38]. Unfortunately, real world networks are characterized by a high clustering coefficient [18], so often the analytical results obtained this way differ quantitatively from ones given by Monte Carlo simulations [37]. The need to sweep a multidimensional space of input parameters and perform multiple independent repetitions to obtain good statistical results makes Monte Carlo simulations very time-consuming. For this reason, we utilize both computer simulations and analytical methods in this research.
@@
A popular agent-based model that already has found its use in studying diffusion of eco-innovation is the $q$-voter model [20, 23, 39]. In this model, conformity is the basic form of social response. Agents are characterized by a single binary variable denoting their opinion and placed in nodes of some underlying graph structure. Their opinions change upon impact with so-called groups of influence, i.e. $q$ of the agent’s neighbors chosen randomly, but only if a given group presents an unanimous opinion. A well-established extension to the $q$-voter model is the addition of independence – the probability that an agent acts and changes opinion independently of the group of influence [40]. The $q$-voter model has been already examined on multi-layer networks, but to the best of our knowledge only ones consisting of two identical layers [35, 41].
@@
In this paper, we introduce a new model based on the aforementioned $q$-voter model with independence. Our goal is to capture both opinion dynamics and diffusion of innovation jointly. Hence, we introduce (compared to the original $q$-voter model) a second agent’s attribute, in addition to the opinion – an adoption state. Similarly to the opinion, it is a binary one. Positive value means that an agent possesses a PV installation, negative – it does not. An agent (in this case a household) can acquire knowledge of PVs from two sources. Either it sees panels on the roofs of neighboring households, or communicates with friends/colleagues. To model this twofold dynamic, we utilize a two-layer network as the underlying topology. The first layer of the network depicts the spacial location of an agent, similarly to [42, 43]. From a mathematical point of view it is just a square lattice (SL) with Moore’s neighborhood [44]. At this level, each agent possesses only knowledge of the adoption states of the closest neighborhood, and these adoption states shape agent’s opinion, according to the $q$-voter rule. The agent cannot share opinion or be subject to the opinions of the neighborhood. This is the visual observation part. A single-family house owner sees only PV panels, or lack thereof, on the roofs of neighboring houses, and does not exchange information or opinions with neighbors. The second layer of the network corresponds to structure of contacts and relationships of agents. For this reason, here we use two-dimensional Watts-Strogatz graph (WS2D) [45]. It possesses some characteristics of real world social networks [18, 46] and, in a specific case (when randomness equals 00), reduces to a square lattice. At this level, a connection between two agents implies their familiarity, through which they can exchange opinions, but do not observe their adoption states. In this case, social influence is also described by the $q$-voter model. Lastly, it is an agent’s opinion solely that impacts its adoption state. An agent with a positive opinion can install photovoltaic panels with probability $a_{1}$, one with a negative one can uninstall with probability $a_{2}$. The reason for this research is to thoroughly examine this mathematical model, to understand how its parameters shape the outcome.
@@
The remainder of this article is organized as follows. In the next section, we provide a more detailed description of the model and the methods used. Then, in Section 3, we examine the two variants, AND and OR rules, which were explained above. Both by means of computer simulations, numerical and analytical methods. Finally, in Section 4, we wrap up. Additionally, in the Appendix, we consider various initial conditions.
@@ 2.1 Simulation Model
We consider a set of $N$ agents, each of which is characterized by 2 binary variables: adoption $A_{i}=\pm 1$ (adopted or not adopted), and opinion $S_{i}=\pm 1$ (positive or negative), for $i=1,2,\dots,N$. Agents are located on a two-layer network (Fig. 2). The first layer of the network depicts the spacial location of an agent and is represented by a square lattice (SL) with Moore’s neighborhood, meaning that each agent possesses 8 neighbors surrounding it (except for agents in the corners and along the edges, which have 3 and 5 neighbors, respectively) [44]. At this level, an agent sees only the adoption states of its neighbors. The second layer corresponds to the structure of social ties of agents, which is described by a two-dimensional Watts-Strogatz graph [45]. Here, an agent sees only the opinions of its neighbors. Each layer is a connected simple graph [47], as described in Definition 1.
@@
Let $G=\left[G_{X,i,j}\right]$ denote the adjacency matrix for the two-layer network, with each layer $X$, $X\in\{1,2\}$, being a connected simple graph, i.e. an unweighted, undirected graph containing no graph loops or multiple edges [47]. Then, for $i,j=1,2,\dots,N$:
@@
We investigate the model through Monte Carlo simulations with a random sequential updating scheme. Within a single simulation, the time is measured in so-called Monte Carlo steps (MCS) [48]. One MCS consists of $N$ elementary events, each of length $\Delta t=\frac{1}{N}$. In each elementary event, an agent $i$ is chosen at random (uniformly from all $N$). Then, with probability $p$, agent $i$ acts independently and changes its opinion $S_{i}$ randomly. With complementary probability $1-p$, the agent is susceptible to social influence (based on the $q$-voter model [39]), combined from both layers of the network with respect to the variant, AND or OR. Lastly, if agent $i$ has positive opinion $S_{i}=+1$, but negative adoption state $A_{i}=-1$, it changes the latter to positive, $A_{i}\to+1$, with probability $a_{1}$. Otherwise, if it possesses negative opinion $S_{i}=-1$, but positive adoption state $A_{i}=+1$, it looses the latter, i.e. $A_{i}\to-1$, with probability $a_{2}$. Simulation details are shown in Algorithm 1, and the graphical representation of the model in Fig. 2. In Algorithm 1:
@@
    $G$ is the adjacency matrix for the two-layer network, as per Definition 1,
@@
To clarify, we do the following. At the beginning of each independent simulation (trajectory), we set the initial conditions. In simulations, we always start from a fully unadopted, negative state, i.e., $\forall_{i}\,\,A_{i}(0)=-1,\,\,S_{i}(0)=-1$. Then, we perform Monte Carlo steps until a determined time horizon $T$ is reached ($T$ is the number of MCS). We repeat independent simulations multiple times with the same set of parameter values for a better statistical accuracy. We also run separate simulations for different sets of parameter values.
@@
$c_{A}=\frac{1}{2N}\sum_{i=1}^{N}\left(A_{i}+1\right),\,\,\,\,c_{S}=\frac{1}{2N}\sum_{i=1}^{N}\left(S_{i}+1\right).$ (2)
@@ 2.2 Mean-Field Approximation
Let $c_{A}$ denote the concentration of positive adoption states and the probability of choosing an agent with a positive adoption state, and $c_{S}$ – the concentration of positive opinions and the probability of choosing an agent with a positive opinion. Under the assumptions that these events are independent, and that each layer of the network is of size $N\to\infty$, the dynamics of the system is described in the AND variant by:
@@
Under our assumptions, we can write down probability $\gamma_{A}^{+}$ that the number of agents with positive adoption states will increase by $1$ and probability $\gamma_{A}^{-}$ that it will decrease by $1$:
@@
which gives us Eq. (3), after replacing $a_{2}$ with $ha_{1}$, see Eq. (1). Eq. (3) is common for both AND and OR variants of the model. Similarly, we can write down probabilities $\gamma_{S}^{+}$ and $\gamma_{S}^{-}$ that a number of agents with positive opinion will increase or decrease by $1$, respectively. These are:
@@
Due to the power of $q$ in Eqs. (4)-(5), time trajectories of the system given by Eqs. (3)-(5) cannot be determined analytically. Hence, one must resort to numerical methods, which we show in Section 3. However, stationary states can be found analytically and their existence can be proven.
@@
Let $c_{A}$ denote the concentration of positive adoption states and the probability of choosing an agent with a positive adoption state, and $c_{S}$ – the concentration of positive opinions and the probability of choosing an agent with a positive opinion. Under the assumptions that these events are independent, and that each layer of the network is of size $N\to\infty$, for any $q\in\mathbb{N}$, $q\geq 2$, $p\in[0,1]$, $h\in(0,1)$ and $a_{1}\in(0,1]$, there always exists at least one stationary state, and all stationary states must satisfy:
@@
A stationary state must satisfy:
@@
When $c_{S}\in\left[\frac{1}{2},1\right]$, $f_{4}(c_{S})$ is a rational function with a positive denominator, and hence continuous within that interval. As it is continuous in $\left[\frac{1}{2},1\right]$, $f_{4}(1)=0$ and $f_{4}\left(\frac{1}{2}\right)=1$, it achieves any value between 00 and $1$. Hence, for any $p\in[0,1]$ there always exists at least one $c_{S}\in[0,1]$ that satisfies Eq. (30). Therefore, in the AND variant, there always exists at least one stationary state.
@@
where $f_{3}(c_{S})$ is given by Eq. (27) and
@@
When $c_{S}\in\left[\frac{1}{2},1\right]$, $f_{6}(c_{S})$ is a rational function with positive denominator, and hence continuous within that interval. As it is continuous in $\left[\frac{1}{2},1\right]$, $f_{6}(1)=0$ and $f_{6}\left(\frac{1}{2}\right)=1$, it achieves any value between 00 and $1$. Hence, for any $p\in[0,1]$ there always exists at least one $c_{S}\in[0,1]$ that satisfies Eq. (36). Therefore, in the OR variant, there always exists at least one stationary state. ∎
@@
First, we note that for $c_{S}\in\left[\frac{1}{2},1\right)$:
@@ 3.1 Simulation Details
First, we examine our model with Monte Carlo computer simulations. In order to measure the outcomes at the macroscopic level, we use concentrations, i.e. fractions, of positive adoption states $c_{A}$ and positive opinions $c_{S}$, see Definition 2. In all the simulations, we study a system of size $N=2500$ (number of agents). As mentioned in the Introduction, the first layer of the two-layer network is always a Square Lattice with Moore’s neighborhood [44], and no periodic boundary conditions, which means that almost every agent has exactly 8 neighbors (except those on the edges of the lattice). We denote it by SL($N=2500$, $m=1$), where $m=1$ indicates the range of a neighborhood on a square grid. The second layer is always a two-dimensional Watts-Strogatz graph [45], i.e. a Square Lattice with Moore’s neighborhood before rewiring and rewiring probability $\beta=0.2$, and again, without periodic boundary conditions. We denote it by WS2D($N=2500$, $m=1$, $\beta=0.2$). We study the behavior of the system with respect to the parameters $p$, $a_{1}$ and $a_{2}$. Although the size of the group of influence $q$ is a parameter as well, here we keep it constant, as its impact in the $q$-voter model with independence is well studied [40]. We choose $q=4$ motivated by real life experiments [50]. This will help us to reduce computational efforts. In all the simulations, we start from a fully unadopted, negative state, i.e., $c_{A}(0)=0$, $c_{S}(0)=0$. Lastly, for illustrative purposes, the time horizon $T$ for each simulation is limited to $5000$ Monte Carlo steps (MCS), and the number of time trajectories to $10$.
@@ 3.2 Simulation Model
Figure 3: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Let us begin with Fig. 3. It shows sample evolutions of the system, for both variants (AND and OR), and a range of values of $p$, $a_{1}$ and $a_{2}$, where $a_{2}=0.5a_{1}$ in all cases. We can observe that several final outcomes are possible. When $p$ is low, we retain an unadopted state ($c_{A}(T)\approx 0$, $c_{S}(T)\approx 0$). None of the agents have a positive opinion or adoption state, except for a few “rebels”. Then, there is a critical value of $p$, above which the system becomes adopted ($c_{A}(T)\approx 1$, $c_{S}(T)\approx 1$). If the value of $p$ is high, independence surpasses conformity and we end up with a disordered system ($c_{A}(T)\approx 0.5$, $c_{S}(T)\approx 0.5$). There is a visible difference between the two variants: AND and OR. In the first, much lower values of $p$ are required to enter adopted or disordered state than in the latter. Finally, the value of $a_{1}$ itself has no impact on the final state (except for $a_{1}=0$, which is not shown here), only on the time needed to reach it, with the higher value speeding up the process.
Figure 4: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for different values of $p$: the AND (top), the OR variant (bottom). Values of $a_{1}$: $a_{1}=0.04$ (left), $a_{1}=0.16$ (right). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.25a_{1}$ in all cases.
What affects the final state, however, is the relationship between $a_{1}$ and $a_{2}$, namely the coefficient $h$, see Eq. (1). In Fig. 4, $a_{2}$ is twice smaller than before, i.e. $a_{2}=0.25a_{1}$. With this decrease in $a_{2}$ much lower values of $p$ are needed, in both variants, for the system to become adopted. This itself is a rather trivial conclusion, as now the probability of loosing adoption is four times smaller that the one of adopting. There is, however, a secondary effect to it, due to the fact how the first layer of the network impacts opinions in our model. Positive adoption states support positive opinions (and vice versa). For this reason, the adopted state is not only easier to reach, but also more difficult to disorder.